# MetaSym: A Symplectic Meta-learning Framework for Physical Intelligence

### v2

@@ 3.2 ActiveDecoder: Autoregressive Decoder
As mentioned in Section 1.1, we train an autoregressive decoder to model non conservative and realistic forces such friction, or air resistance, that break pure Hamiltonian symmetries. Unlike the SymplecticEncoder, which strictly enforces canonical updates, the ActiveDecoder can incorporate these extraneous phenomena.
We define $\mathbf{\tilde{q}^{(i)}_{t}}$ and $\mathbf{\tilde{p}^{(i)}_{t}}$ as the dissipative canonical coordinates, of a system $i$. An input sequence $\left(\mathbf{\tilde{q}^{(i)}_{t:c}},\mathbf{\tilde{p}^{(i)}_{t:c}},\Delta t,\mathbf{u^{(i)}_{t:c}}\right)\in\mathbb{R}^{(2d+m+1)\times c}$ with a context-window of c timesteps is provided to a linear projection, where $\mathbf{u^{(i)}_{t}}\in\mathbb{R}^{m}$ represents the m-dimensional control-input driving the system at time $t$. This linear layer, which we call “ControlNet”, serves to map the input to a latent vector $\mathbf{z}_{d}\in\mathbb{R}^{2d}$ representing the non-conservative part of the system’s dynamics. We then apply a masked multi-head self-attention over the sequence $\mathbf{z}_{d}$ for autoregressive decoding. The masking allows us to model the causal dependencies of the decoder’s input sequence.
Next, we apply a cross-attention mechanism, augmented with a specialized meta-attention design. To achieve multi-system generalization and fast adaptation, we use a meta-learning setup akin to a bi-level optimization inspired by recent progress [31]. We separate the ActiveDecoder’s parameters into: global parameters, $\theta_{AD}$, that remain shared across all systems, and local parameters, $\zeta_{i}$, that can be interpreted as system specific parameters. More specifically, in the meta-attention:
@@
For each system $i$, we split the time-sequenced data, $\mathbf{\tilde{q}^{(i)}},\mathbf{\tilde{p}^{(i)}},\mathbf{u^{(i)}}$ into an adaptation set , $\mathcal{I}_{adapt}$ containing $\mathcal{T}_{adapt}$ number of timesteps, and a meta set $\mathcal{I}_{meta}$ with $\mathcal{T}_{meta}$ number of timesteps. During the inner loop of the bi-level optimization, we hold $\theta_{AD}$ fixed and fine-tune $\zeta_{i}$ by performing a few gradient steps. Each step iterates over $\mathcal{I}_{adapt}$ feeding the ActiveDecoder ($\Phi_{AD}$) with ground-truth $(\mathbf{q}_{t:c}^{(i)},\mathbf{p}_{t:c}^{(i)})$ and control $\mathbf{u}_{t:c}^{(i)}$, and minimizing an inner MSE loss across $\mathcal{T}_{adapt}$ number of timesteps:
$\mathcal{L}_{\mathrm{inner}}^{(i)}=\frac{1}{\mathcal{T}_{adapt}}\sum_{t\in\mathcal{I}_{\mathrm{adapt}}}\Bigl{\|}\Phi_{\theta_{\text{AD}}}$ $\bigl{(}\mathbf{q}_{t:c}^{(i)},\mathbf{p}_{t:c}^{(i)},\mathbf{u}_{t:c}^{(i)};\boldsymbol{\theta_{AD}},\boldsymbol{\zeta}_{i}\bigr{)}-[\mathbf{q}_{t+1}^{(i)},\mathbf{p}_{t+1}^{(i)}]\Bigr{\|}^{2}.$ (4)
@@
Once the local parameters $\zeta_{i}$ have been adapted to the optimal value $\zeta^{*}_{i}$, we evaluate the model’s performance on $\mathcal{I}_{meta}$ portion of the trajectory. The corresponding outer loss:
$\mathcal{L}_{\mathrm{outer}}^{(i)}=\frac{1}{\mathcal{T}_{meta}}\sum_{t\in\mathcal{I}_{\mathrm{meta}}}\Bigl{\|}\Phi_{\theta_{\text{AD}}}$ $\bigl{(}\mathbf{q}_{t:c}^{(i)},\mathbf{p}_{t:c}^{(i)},\mathbf{u}_{t:c}^{(i)};\boldsymbol{\theta_{AD}},\boldsymbol{\zeta}^{*}_{i}\bigr{)}-[\mathbf{q}_{t+1}^{(i)},\mathbf{p}_{t+1}^{(i)}]\Bigr{\|}^{2},$ (5)
propagates gradients only to $\theta_{AD}$. The overall training objective is then the sum over $\mathcal{L}_{\mathrm{outer}}^{(i)}$ for all systems $i$, over the batch. We finalize the decoder with a Multi-Layer Perceptron (MLP) that outputs the position and momentum for the next timestep $\left(\mathbf{\hat{q}^{(i)}_{t+1}},\mathbf{\hat{p}^{(i)}_{t+1}}\right)$. The architecture is then used autoregressively for future predictions. The inherent high dimensionality of the ActiveDecoder’s internal representations, coupled with the lack of well-defined or tractable Hessian guarantees in this setting [32], motivates our adoption of this meta-learning paradigm over traditional methods such as MAML. By selectively adapting only a carefully chosen subset of ActiveDecoder’s parameters, we are able to minimize both the number of adaptation steps and the quantity of data required during the meta-learning process.
@@ 3.3 Autoregressive Inference
Once both the SymplecticEncoder and ActiveDecoder (Sections 3.1 and 3.2) are trained, we generate future predictions by rolling out the model autoregressively during inference time. Specifically, at test time, we are given initial phase-space measurements. We first map to the conservative portion of the phase-space, ($\mathbf{q}^{(i)}_{t},\mathbf{p}^{(i)}_{t}$) through the SymplecticEncoder. The ActiveDecoder then produces the next predicted phase-space trajectory $(\mathbf{\hat{q}^{(i)}_{t+1}},\mathbf{\hat{p}^{(i)}_{t+1}})$, based on the pipeline highlighted in Section 3.2. However, we subsequently treat $(\mathbf{\hat{q}^{(i)}_{t+1}},\mathbf{\hat{p}^{(i)}_{t+1}})$ as inputs for the next timesteps. This is repeated over the context window provided to the decoder.
In summary, by pairing our SymplecticEncoder (to ensure structure preservation) that provides a strong inductive bias with an Autoregressive Transformer-style Decoder (ActiveDecoder) equipped with meta-attention (to handle individual system variations), MetaSym can rapidly adapt to diverse physical systems while guaranteeing physically consistent core dynamics. The proof of this strong inductive bias provided by the SymplecticEncoder and its effect on the training of the ActiveDecoder as part of MetaSym is provided by Appendix C. The sketch of this proof is as follows, the SymplecticEncoder $\Phi_{\theta_{\mathrm{SE}}}$ exactly preserves the form $d\Phi_{\theta_{\mathrm{SE}}}(\mathbf{x})^{\top}J\,d\Phi_{\theta_{\mathrm{SE}}}(\mathbf{x})=J$, while the ActiveDecoder can be written as $\Phi_{\theta_{\mathrm{AD}}}(\mathbf{z}_{c},\mathbf{u}_{t},\mathbf{d}_{t})$ where $\mathbf{u}_{t}$ and $\mathbf{d}_{t}$ represent the seperable, for the sake of this proof, control-input and dissipation. The Jacobian $\partial\Phi_{\theta_{\mathrm{AD}}}=I+A$ and $\|A\|\leq\rho$, where $\rho$ is shown empirically be bounded even under extreme deformation. Composing these maps gives $\Phi=\Phi_{\theta_{\mathrm{AD}}}\circ\Phi_{\theta_{\mathrm{SE}}}$ and hence $d\Phi=(I+A)\,d\Phi_{\theta_{\mathrm{SE}}}$. Expanding the symplectic form then yields, $d\Phi^{\top}J\,d\Phi=d\Phi_{\theta_{\mathrm{SE}}}^{\top}\bigl{(}J+A^{\top}J+JA+A^{\top}JA\bigr{)}\,d\Phi_{\theta_{\mathrm{SE}}}=J+O(\rho),$ which shows that $\Phi$ deviates from exact symplecticity by at most $O(\rho)$. An identical $O(\rho)$ argument on $\det d\Phi=\det(I+A)\,\det d\Phi_{\mathrm{SE}}$ establishes near–volume preservation, completing the proof that $\Phi$ is $O(\rho)$–near-symplectic.
@@ 4 Results
In this section, we evaluate MetaSym’s performance, against a variety of State-of-The-Art (SoTA) models and benchmarks, especially in long-horizon stability and few-shot generalization. In particular, we report the behavior of MetaSym against Transformers that have achieved impressive performance in modeling physical systems [19] and DHNNs [17] that model systems via the Hemholtz decomposition, although they require gradient information and a numerical integrator to obtain trajectory data. Our model predicts the subsequent timesteps directly, which poses a much harder task that does not require numerical integration, which can be time-consuming and may hinder real-time performance. These results also reflect on the effectiveness of our design choices.
@@
As mentioned in Section 1.1, we choose to benchmark MetaSym for three extremely challenging and diverse datasets. Notably, we choose the spring mesh system, which tests scalability and closely resembles finite element modeling of surfaces with different materials. This dataset has provided the state-of-art testbed for large dimensional systems [1]. Moreover, we utilize a well proven, reproducible dataset of an open-quantum system under undergoing heterodyne detection leading to representative quantum effects such as decoherence to demonstrate MetaSym’s robustness to noise and flexibility to model all types of physical systems. Finally, we choose to predict the dynamics of a quadrotor, a long-standing benchmark in robotics and challenging due to floating-base dynamics and sensor noise that is simulated using standard Gaussian additive noise [33]. The description of each experimental setup is laid in the next sections, however a more extensive description can be found in Appendix B. Finally, information regarding the utilized compute resources and training hyperparameters are provided in Appendices E & F.
@@
Figure 2: (Left) Time evolution of the system’s position and momentum variables for a representative set of masses in the spring mesh. The orange curves represent the ground-truth trajectories for each phase-space coordinate type (i.e., $q_{x},q_{y},p_{x}$, and $p_{y}$), while the blue curves depict the model’s predictions. The close alignment between these trajectories underscores the model’s capacity to accurately capture the underlying long-term dynamics (600 timesteps) of the coupled spring system using a context window of 30 timesteps. (Right) Plots illustrating the mean squared error (MSE) of the time evolution of each phase-space coordinate type (dots) for five (A-E) spring-mesh systems in the test set. Each column encapsulates the spread of errors observed across all masses in the spring-mesh for a given phase-space coordinate across multiple timesteps, with the boxes marking their median values. The uniformly low median errors across all components demonstrate that the model generalizes effectively to different spring-mesh systems for all phase-space coordinates.
@@ 4.1 Spring-Mesh System
As our first benchmark, we use the 400-dimensional spring-mesh system introduced by Otness et al. [1], consisting of a $10\times 10$ lattice of point masses connected by elastic springs. Each of the 100 masses possesses two positional and two momentum degrees of freedom, yielding a 400-dimensional state space. This system, notable for its high dimensionality and complex dynamical couplings, serves as a canonical testbed for learning physical systems, including deformable volumes and surfaces relevant to engineering applications [34]. Node positions are updated via forces from spring extensions and compressions, resulting in nontrivial communication and computation patterns across the mesh. For a more in-depth description of the benchmark, refer to Appendix B.1. Unless stated otherwise, models are trained using teacher forcing for next-token prediction with a fixed context window of 30 timesteps. This choice aligns with prior configurations used in large-scale visual-language architectures (e.g., $\pi 0$: 50 timesteps [35], OpenVLA: 8 timesteps [36]). Notably, our models are significantly smaller, yet demonstrate strong capability for long-term, temporally consistent predictions within this horizon as the results in Figure 2 indicate.
Figure 3: (Left) Time evolution of the two quadratures measured via heterodyne detection for a representative quantum system in the test set, characterized by a measurement efficiency $\eta=0.86$ and measurement strength $\kappa$. The orange lines indicate the true measurement trajectories, while the blue lines display the model’s predictions (MetaSym). The close overlap between these trajectories highlights the model’s effectiveness in capturing the underlying quantum dynamics from heterodyne signals. The context window of the model is 30 timesteps. (Right) Plots showing the mean squared error (MSE) of the predicted quadratures (dots) for five randomly selected test systems. The consistently low median errors (boxes) across all tested systems underscore the robustness and generalization capabilities of the model.
@@ 4.2 Open Quantum System
To benchmark MetaSym on a novel open quantum dynamics task, we consider a parametric oscillator initialized in a coherent state [37]. The Hamiltonian includes a harmonic term $\hat{H}{{}_{\mathrm{osc}}}=\omega\hat{a}^{\dagger}\hat{a}$, a squeezing term $\hat{H}{{}_{\mathrm{sqz}}}=\frac{i\chi}{2}(\hat{a}^{\dagger 2}-\hat{a}^{2})$, and a cubic drive $\hat{H}{{}_{\text{cubic}}}=\beta(\hat{a}^{3}+\hat{a}^{\dagger 3})$, with $\hat{a},\hat{a}^{\dagger}$ the ladder operators in a truncated Fock space of dimension $N$. Dissipation arises via coupling to a thermal bath with rate $\gamma$ and mean occupation $\tilde{n}{{}_{\mathrm{th}}}$. Continuous heterodyne detection yields a stochastic master equation for the conditional state (Appendix B.2). The only experimentally accessible data are the real and imaginary quadratures $X$ and $P$ of the heterodyne current, extracted from single-shot trajectories, as the quantum state itself remains inaccessible. We simulate this setup by numerically solving the stochastic master equation [38], varying measurement efficiency $\eta$, squeezing strength $\chi$, cubic drive $\beta$, oscillator frequency $\omega$, bath occupation $\tilde{n}{{}_{\mathrm{th}}}$, and initial states. Training and out-of-distribution inference details are in Appendix B.2. As shown in Figure 3, MetaSym performs well in this strongly stochastic, non-classical regime.
Figure 4: (Left) Represents the translational and angular phase-space evolution of the quadrotor, with training and test trajectories generated using the Crocoddyl trajectory optimization package [39]. Each task is initialized with randomized initial conditions and a randomized terminal position over a 1.5s horizon (150 timesteps) with a context window of 30 timesteps. The ground-truth test trajectory (orange line) overlaps with MetaSym’s predictions (blue line) indicating the excellent predictive capabilites of our model. (Right) Plots summarizing the mean squared error (MSE) of the phase-space coordinates evolution for five randomly generated test systems (dots). The consistently low median errors (boxes) across all components underscore the robustness and generalization capabilities of the model.

### v3

@@ 3.2 ActiveDecoder: Autoregressive Decoder
As mentioned in Section 1.1, we train an autoregressive decoder to model non-conservative and realistic forces such friction, or air resistance, that break pure Hamiltonian symmetries. Unlike the SymplecticEncoder, which strictly enforces canonical updates, the ActiveDecoder can incorporate these extraneous phenomena.
We define $\mathbf{\tilde{q}^{(i)}_{t}}$ and $\mathbf{\tilde{p}^{(i)}_{t}}$ as the dissipative canonical coordinates, of a system $i$. An input sequence $\left(\mathbf{\tilde{q}^{(i)}_{t:c}},\mathbf{\tilde{p}^{(i)}_{t:c}},\Delta t,\mathbf{u^{(i)}_{t:c}}\right)\in\mathbb{R}^{(2d+m+1)\times c}$ with a context-window of c time-steps is provided to a linear projection, where $\mathbf{u^{(i)}_{t}}\in\mathbb{R}^{m}$ represents the $m$-dimensional control-input driving the system at time $t$. This linear layer, which we call ControlNet as part of the ActiveDecoder serves to map the input to a latent vector $\mathbf{z}_{d}\in\mathbb{R}^{2d}$ representing the non-conservative part of the system’s dynamics. We then apply a masked multi-head self-attention over the sequence $\mathbf{z}_{d}$ for autoregressive decoding. The masking allows us to model the causal dependencies of the decoder’s input sequence.
Next, we apply a cross-attention mechanism, augmented with a specialized meta-attention design. To achieve multi-system generalization and fast adaptation, we use a meta-learning setup akin to a bi-level optimization inspired by recent progress (Li et al., 2025). We separate the ActiveDecoder’s parameters into global parameters, $\theta_{AD}$ that remain shared across all systems, and local parameters, $\zeta_{i}$ that can be interpreted as system specific parameters. More specifically, in the meta-attention:
@@
For each system $i$, we split the time-sequenced data, $\mathbf{\tilde{q}^{(i)}},\mathbf{\tilde{p}^{(i)}},\mathbf{u^{(i)}}$ into an adaptation set, $\mathcal{I}_{adapt}$ containing $\mathcal{T}_{adapt}$ number of time-steps, and a meta set $\mathcal{I}_{meta}$ with $\mathcal{T}_{meta}$ number of time-steps. During the inner loop of the bi-level optimization, we hold $\theta_{AD}$ fixed and fine-tune $\zeta_{i}$ by performing a few gradient steps. Each step iterates over $\mathcal{I}_{adapt}$ feeding the ActiveDecoder ($\Phi_{AD}$) with ground-truth $(\mathbf{q}_{t:c}^{(i)},\mathbf{p}_{t:c}^{(i)})$ and control $\mathbf{u}_{t:c}^{(i)}$, and minimizing an inner MSE loss across $\mathcal{T}_{adapt}$ number of time-steps, i.e.,
$\mathcal{L}_{\mathrm{inner}}^{(i)}=\frac{1}{\mathcal{T}_{adapt}}\sum_{t\in\mathcal{I}_{\mathrm{adapt}}}\Bigl\|\Phi_{\theta_{\text{AD}}}\bigl(\mathbf{q}_{t:c}^{(i)},\mathbf{p}_{t:c}^{(i)},\mathbf{u}_{t:c}^{(i)};\bm{\theta_{AD}},\bm{\zeta}_{i}\bigr)-[\mathbf{q}_{t+1}^{(i)},\mathbf{p}_{t+1}^{(i)}]\Bigr\|^{2}.$ (5)
@@
Once the local parameters $\zeta_{i}$ have been adapted to the optimal value $\zeta^{*}_{i}$, we evaluate the model’s performance on $\mathcal{I}_{meta}$ portion of the trajectory. The corresponding outer loss,
$\mathcal{L}_{\mathrm{outer}}^{(i)}=\frac{1}{\mathcal{T}_{meta}}\sum_{t\in\mathcal{I}_{\mathrm{meta}}}\Bigl\|\Phi_{\theta_{\text{AD}}}\bigl(\mathbf{q}_{t:c}^{(i)},\mathbf{p}_{t:c}^{(i)},\mathbf{u}_{t:c}^{(i)};\bm{\theta_{AD}},\bm{\zeta}^{*}_{i}\bigr)-[\mathbf{q}_{t+1}^{(i)},\mathbf{p}_{t+1}^{(i)}]\Bigr\|^{2},$ (6)
propagates gradients only to $\theta_{AD}$. The overall training objective is then the sum over $\mathcal{L}_{\mathrm{outer}}^{(i)}$ for all systems $i$, over the batch. We finalize the decoder with a Multi-Layer Perceptron (MLP) that outputs the position and momentum for the next time-step $\left(\mathbf{\hat{q}^{(i)}_{t+1}},\mathbf{\hat{p}^{(i)}_{t+1}}\right)$. The architecture is then used autoregressively for future predictions. The inherent high dimensionality of the ActiveDecoder’s internal representations, coupled with the lack of well-defined or tractable Hessian guarantees in this setting (LeCun et al., 2012), motivates our adoption of this meta-learning paradigm over traditional methods such as MAML. By selectively adapting only a carefully chosen subset of ActiveDecoder’s parameters, we are able to minimize both the number of adaptation steps and the quantity of data required during the meta-learning process.
@@ 3.3 Autoregressive Inference
[…]
In summary, by pairing our SymplecticEncoder (to ensure structure preservation) that provides a strong inductive bias with an Autoregressive Transformer-style Decoder (ActiveDecoder) equipped with meta-attention (to handle individual system variations), MetaSym can rapidly adapt to diverse physical systems while guaranteeing physically consistent core dynamics. The proof of this strong inductive bias provided by the SymplecticEncoder and its effect on the training of the ActiveDecoder as part of MetaSym is provided by Appendix C.
@@ 4 Results
In this section, we evaluate MetaSym’s performance, against a variety of SOTA models and benchmarks, especially in long-horizon stability and few-shot generalization. In particular, we report the behavior of MetaSym against Transformers that have achieved impressive performance in modeling physical systems (Geneva & Zabaras, 2022) and DHNNs (Sosanya & Greydanus, 2022) that model systems via the Helmholtz decomposition, although they require gradient information and a numerical integrator to obtain trajectory data. Our model predicts the subsequent time-steps directly, which poses a much harder task that does not require numerical integration, which can be time-consuming and may hinder real-time performance. These results also reflect on the effectiveness of our design choices.
@@
As mentioned in Section 1.1, we choose to benchmark MetaSym for three extremely challenging and diverse datasets. Notably, we choose the spring-mesh system, which tests scalability and closely resembles finite element modeling of surfaces with different materials. This dataset has provided the SOTA testbed for large dimensional systems (Otness et al., 2021). Moreover, we utilize a well proven, reproducible dataset of an open-quantum system undergoing heterodyne detection leading to representative quantum effects, such as decoherence, to demonstrate MetaSym’s robustness to noise and flexibility to model all types of physical systems. Finally, we choose to predict the dynamics of a quadrotor, a long-standing benchmark in robotics and challenging due to floating-base dynamics and sensor noise that is simulated using standard Gaussian additive noise (Bansal et al., 2016). Furthermore, we also demonstrate the adaptability of MetaSym to real-world data by adapting this model (Mohajerin & Waslander, 2018). The description of each experimental setup is laid in the next sections, however a more extensive description can be found in Appendix B. All reported MSE errors are not weighted. In the plots, the errors correspond to each phase-space coordinate, while in Table 1 the reported errors are calculated across all phase-space dimensions. Finally, information regarding the utilized compute resources and training hyperparameters are provided in Appendices F & G.
@@
Figure 2: (Left) Time evolution of the system’s position and momentum variables for a representative set of masses in the spring mesh. The orange curves represent the ground-truth trajectories for each phase-space coordinate type (i.e., $q_{x},q_{y},p_{x}$, and $p_{y}$), while the blue curves depict the model’s predictions. The close alignment between these trajectories underscores the model’s capacity to accurately capture the underlying long-term dynamics (600 time-steps) of the coupled spring system using a context window of 30 time-steps. (Right) Plots illustrating the mean squared error (MSE) of the time evolution of each phase-space coordinate type (dots) for five (A-E) spring-mesh systems in the test set. Each column encapsulates the spread of errors observed across all masses in the spring-mesh for a given phase-space coordinate across multiple time-steps, with the boxes marking their median values. The uniformly low median errors across all components demonstrate that the model generalizes effectively to different spring-mesh systems for all phase-space coordinates.
@@ 4.1 Spring-mesh System
As our first benchmark, we use the 400-dimensional spring-mesh system introduced by Otness et al. (2021), consisting of a $10\times 10$ lattice of point masses connected by elastic springs. Each of the 100 masses possesses two positional and two momentum degrees of freedom, yielding a 400-dimensional state space. This system, notable for its high dimensionality and complex dynamical couplings, serves as a canonical testbed for learning physical systems, including deformable volumes and surfaces relevant to engineering applications (Pfaff et al., 2021). Node positions are updated via forces from spring extensions and compressions, resulting in nontrivial communication and computation patterns across the mesh. For a more in-depth description of the benchmark, refer to Appendix B.1. Unless stated otherwise, models are trained using teacher forcing for next-token prediction with a fixed context window of 30 time-steps. This choice aligns with prior configurations used in large-scale visual-language architectures (e.g., $\pi 0$: 50 time-steps in Black et al. (2024), OpenVLA: 8 time-steps in Kim et al. (2025)). We also provide an ablation study on the context window, refer to Appendix D. Notably, our models are significantly smaller, yet demonstrate strong capability for long-term, temporally consistent predictions within this horizon as the results in Fig. 2 indicate.
Figure 3: (Top) Time evolution of the two quadratures measured via heterodyne detection for a representative quantum system in the test set, characterized by a measurement efficiency ($\eta=0.86$) and measurement strength ($\kappa$). The orange lines indicate the true measurement trajectories, while the blue lines display the model’s predictions (MetaSym). The close overlap between these trajectories highlights the model’s effectiveness in capturing the underlying quantum dynamics from heterodyne signals. The context window of the model is 30 time-steps. (Bottom) Plots showing the mean squared error (MSE) of each of the predicted quadratures (dots) for five randomly selected test systems. The consistently low median errors (boxes) across all tested systems underscore the robustness and generalization capabilities of the model.
@@ 4.2 Open Quantum System
To benchmark MetaSym on a novel open quantum dynamics task, we consider a parametric oscillator initialized in a coherent state (Gardiner & Zoller, 2010). The Hamiltonian includes a harmonic term $\hat{H}{{}_{\mathrm{osc}}}=\omega\hat{a}^{\dagger}\hat{a}$, a squeezing term $\hat{H}{{}_{\mathrm{sqz}}}=\frac{i\chi}{2}(\hat{a}^{\dagger 2}-\hat{a}^{2})$, and a cubic drive $\hat{H}{{}_{\text{cubic}}}=\beta(\hat{a}^{3}+\hat{a}^{\dagger 3})$, with $\hat{a},\hat{a}^{\dagger}$ the ladder operators in a truncated Fock space of dimension $N$. Dissipation arises via coupling to a thermal bath with rate $\gamma$ and mean occupation $\tilde{n}{{}_{\mathrm{th}}}$. Continuous heterodyne detection yields a stochastic master equation for the conditional state (Appendix B.2). The only experimentally accessible data are the real and imaginary quadratures $X$ and $P$ of the heterodyne current, extracted from single-shot trajectories, as the quantum state itself remains inaccessible. In this context, quadratures ($X$ and $P$) denote the non-commuting observables, serving as the quantum optical analogues to classical position and momentum (Milburn, 1987). We simulate this setup by numerically solving the stochastic master equation (Johansson et al., 2012), varying measurement efficiency $\eta$, squeezing strength $\chi$, cubic drive $\beta$, oscillator frequency $\omega$, bath occupation $\tilde{n}{{}_{\mathrm{th}}}$, and initial states. Training and out-of-distribution inference details are in Appendix B.2. As shown in Fig. 3, MetaSym performs well in this strongly stochastic, non-classical regime.
Figure 4: Represents the translational and angular phase-space evolution of the quadrotor, with training generated using the Crocoddyl trajectory optimization package (Mastalli et al., 2020). Each training system is initialized with randomized initial conditions and a randomized terminal position over a $1.5s$ horizon (150 time-steps) with a context window of 30 time-steps. The displayed test trajectories represent the evolution of a real-world quadrotor. The ground-truth test trajectory (orange line) overlaps with MetaSym’s predictions (blue line) indicating the excellent predictive capabilities of our architecture. The MSE across $54$ real-world test trajectories is $\mathbf{0.028\pm 0.091}$ for a context-window of $30$ time-steps. Note we translate the quaternions to roll-pitch-yaw angles to have better interpretability of the results.